40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 13 03:18:14 2011, kiwamu, Update, LSC, Y arm locking prep image.png
    Reply  Tue Sep 13 19:31:09 2011, kiwamu, Update, LSC, Y arm locking prep Yarmlocked.png
       Reply  Wed Sep 14 01:17:51 2011, kiwamu, Update, LSC, measurement of spot position on Y arm misposition.png
          Reply  Wed Sep 14 15:00:00 2011, kiwamu, Update, LSC, measurement of spot position on Y arm DSC_3408_small.jpg
    Reply  Wed Sep 14 21:18:45 2011, kiwamu, Update, LSC, Y arm locking prep 
Message ID: 5392     Entry time: Tue Sep 13 03:18:14 2011     Reply to this: 5398   5410
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: Y arm locking prep 

(Preparation of Y arm locking)

(A) The f2a filters were newly designed and applied to ETMY (see the attachment)

(B) Once the Y arm is aligned such that the TEM00 mode flashes, the transmitted light is visible on the ETMYT CCD camera.

(C) With the newly installed resonant EOM circuit the PDH signal from AS55 looks healthy.

 

(some notes)

(A) To design the f2a filters there is a handy python script called "F2A_LOCKIN.py" in /scripts/SUS.

The script measures the coil imbalance at high frequency and low frequency using a LOCKIN module and then gives us the information about the imbalance.

The script hasn't yet been completed, so it doesn't return the intuitive answers but returns something non-intuitive. I will modify it.

 

(B)  To see the transmitted light from the Y arm I was going to align the CCD camera on the Y end table.

However I found that once the green light is blocked, the transmitted light can be visible on the camera without any re-alignment.

Therefore I haven't rearranged anything on the Y end table, but I just blocked the green light.

Perhaps we still need to align the photo diodes for the transmitted light.

 

(C) While Suresh was working on MC, I looked at the signal from AS55 with all the optics misaligned except for ITMY, ETMY and BS.

The signal from the Y arm looked very PDH signal, and the demodulation phase seemed to be about 45 deg to maximize the I signal.

I tried locking it by feeding the signal back to ETMY but failed due to a too much POS to angle coupling in the ETMY actuators.

I was momentarily able to capture a higher order mode with a negative gain in LSC-YARM_GAIN, but it was quite difficult to keep it locked.

This was because once I increased the gain to make it stable, the angle instability became more significant and lost the lock immediately.

This was the reason why I had to do the f2a filter redesign. Tomorrow we can try locking the Y arm.

Attachment 1: image.png  41 kB  | Hide | Hide all
image.png
ELOG V3.1.3-